A perpentuity due makes annual payments which begins at 100for first year then increses at 6% through the 10th year and then remain level thereafter. Calculate the present value of this perpentuity if the annual effective rate of interest is equal to 8%.
1
Expert's answer
2018-11-20T09:47:10-0500

The present value of this perpentuity if the annual effective rate of interest is equal to 8% is:

PV = D1/(r - g) = 100/(0.08 - 0.06) = 5000.